What is the minimum initial speed of the dart if it is to hit the monkey before the monkey hit the ground, which is 11.2m below the initial position of the monkey, if x=50m and h=10m?(ignore any effects due to air resistance) show the Given, Unknown, and Solution

1
Expert's answer
2022-01-10T09:10:12-0500
"v=x/t=\dfrac{x}{\sqrt{2(11.2-h)/g}}=10.3\text{ m/s}."